# Dissecting users’ needs for search result explanations

### 29 Jan 2024 v1

@@ 6. RQ3: Users’ perception of search explanations and search context provided by search engines
Website previews increase search efficiency: All 12 participants highlighted the usefulness of the website preview feature offered only by Bing that provides an overview of the website’s content. Participants expressed a preference for this feature as it allowed them to get a “whole vibe of the website without visiting it directly” (P2) and “jump to the parts that are relevant to one’s specific search” (P5).
@@
Significance of recommendations along with search results: Bing offers additional search result recommendations alongside the primary search results. Participants found them helpful especially when exploring a topic or seeking specific pieces of information. One participant stated, “I think that’s really helpful when you’re doing something that you’re maybe trying to learn more about a topic and you don’t know where to start… you can kind of follow it down a rabbit hole to something more specific” (P5).
@@
Source information enhances trust: The majority of participants expressed their appreciation for the additional information about the source presented by both Bing and Google, underscoring its role in bolstering the perceived reliability of search results—“ I think linking me to Wikipedia can help me get like a broader understanding of the source with like a decent amount of trust in the results I’m seeing there.” (P9).
@@
Need to improve discoverability of explanation features: For both search engines, participants found it challenging to locate these features, expressing a desire for improved discoverability and accessibility. For example, one participant noted, “I like that it’s there, I wish it was a little bit more transparent so you didn’t have to click on a bubble and like three dots get to that thing” (P5). On similar lines, P11 added, “I mean it’s so deeply layered… I don’t know if I’d ever be able to find it”.
@@ 7.2.1. When explanations are not needed
During the study, when users were not acquainted with the concept of explanations, they revealed that they don’t think too deeply about the search results in their daily lives. Additionally, they only questioned content that appeared irrelevant, inappropriate, or sponsored. This indicates that users generally are not concerned by how search engines curate content as long as their needs are met. Notably, while many participants had desired credibility features from search engines, none mentioned questioning the credibility of search results. This finding is inherently indicative of the amount of trust users have in search results, a phenomenon well-documented in prior literature (Pan et al., 2007).

### 23 Feb 2024 v2

@@ 6.0.5. Website previews increase search efficiency:
All 12 participants highlighted the usefulness of the website preview feature offered only by Bing that provides an overview of the website’s content. Participants expressed a preference for this feature as it allowed them to get a “whole vibe of the website without visiting it directly” (P2) and “jump to the parts that are relevant to one’s specific search” (P5).
6.0.6. Significance of recommendations along with search results:
@@
Bing offers additional search result recommendations alongside the primary search results. Participants found them helpful especially when exploring a topic or seeking specific pieces of information. One participant stated, “I think that’s really helpful when you’re doing something that you’re maybe trying to learn more about a topic and you don’t know where to start… you can kind of follow it down a rabbit hole to something more specific” (P5).
6.0.7. Source information enhances trust:
@@
The majority of participants expressed their appreciation for the additional information about the source presented by both Bing and Google, underscoring its role in bolstering the perceived reliability of search results—“ I think linking me to Wikipedia can help me get like a broader understanding of the source with like a decent amount of trust in the results I’m seeing there.” (P9).
6.0.8. Need to improve discoverability of explanation features:
@@
For both search engines, participants found it challenging to locate these features, expressing a desire for improved discoverability and accessibility. For example, one participant noted, “I like that it’s there, I wish it was a little bit more transparent so you didn’t have to click on a bubble and like three dots get to that thing” (P5). On similar lines, P11 added, “I mean it’s so deeply layered… I don’t know if I’d ever be able to find it”.
@@ 7.1. Users’ needs in evaluating search results
[Google’s topic filters]Figure illustrating Google’s topic filters feature introduced in 2023, showcasing a range of topics associated with the searched topic.
@@ 7.2.1. When explanations are not needed
During the study, when users were not acquainted with the concept of explanations, they revealed that they don’t think too deeply about the search results in their daily lives. Additionally, they only questioned content that appeared irrelevant, inappropriate, or sponsored. This indicates that users generally are not concerned by how search engines curate content as long as their needs are met. Notably, our study also highlighted users trust in the results shown by search engines, a phenomenon well-documented in prior literature (Pan et al., 2007).